Use the following situation to answer the following questions. The table below reports the cost of different quality-relatedWhat is the total cost of preventive activities? (Enter your numerical result only)what is the total cost of fixing problems internally? (i.e., what is the internal failure total costs?) (Enter your numerical
Answer #1

Appraisal costs = Quality audits + Supplier Quality check = 22000 + 4500 = 26500

Preventive costs = Quality planning costs , in this case Special instructions for implementing statistical control = 4500

Total cost of fixing problems internally = Waste (Obsolete material) + Failure cost analysis = 1100 + 8000 = 9100

Note - Return goods technician's salary and Warranty repair costs are external failure costs.